OpenStudy (anonymous):

Soybean meal is 16% protein, cornmeal is 8% protein. How many pounds of each should be mixed together in order to get 320-lb mixture that is 13% protein?

OpenStudy (anonymous):

that question makes almost no sense......

OpenStudy (anonymous):

anyways, just multiply the decimal form of each value by the total weight and that will give you the individual weights for each ingredient

OpenStudy (anonymous):

No im pretty much stumped from this

OpenStudy (anonymous):

very,very,very easy let x=16%= 0.16 let y=8%= 0.08 13%= 0.13 x+y= 320 y= 320 -x 0.16x+0.08y=320*0.13 Substitution : y= 320 -x 0,16x+0.08(320 -x )=320*0.13 here the equation , can u solve from here?

OpenStudy (anonymous):

medal for help?

OpenStudy (anonymous):

%13 can be written as .13 Ask yourself how much protien by mass is in the 360lb mix. 360(.13)=57.6 57.6lbs is your total. The equation can be written (.16)s+(0.08)c=57.6 where s is soybeans and c is cornmeal. Guess and check to determine that 260lbs of soybeans and 200lbs of cornmeal yields 57.6lbs of total protien

OpenStudy (anonymous):

Oops I forgot your total must be 360lbs increase soybean mass nd decrease cornmeal mass...

OpenStudy (anonymous):

0,16x+0.08(320 -x )=320*0.13 0,16x+25.6 -0.08x =41.6 0.08x=41.6-25.6 x=16/0.08 x= 200 Substitution : x= 20 solve for y x+y= 320 200+y= 320 y = 320-200 y = 120 Soybean meal is 200 pounds cornmeal is 120 pounds check; 0.16(200)+0.08(120)=41.6 32+9.6 = 41.6 41.6= 41.6 0.16 or .16 , 0.13 or .13 , 0.08 or .08 is the same 320-lb mixture not 360 lb I don't see that question makes no sense..

OpenStudy (anonymous):

me it make sense

OpenStudy (anonymous):

My answer make sense?